How many dollars do you have if you have 90 nickels, 16 dimes, 16 quarters, and 7 fifty-cent pieces?
Answer by Maths68(1474)

90 nickels = 5*90=450=450/100=$4.5,
16 dimes=16*10=160=160/100=$1.6
16 quarters = 25*16=400=400/100=$4
7 fifty-cent pieces = 7*50=350=350/100=$3.5
Total dollars = 4.5+1.6+4+3.5=$13.6